OpenStudy (anonymous):

i have 3 problems tht are confusing:/ i want ya help simplify: 3x2-15x + 18 ------------ x2-x-6

OpenStudy (callisto):

\[\frac{3x^2-15x + 18}{x^2-x-6}=\frac{3(x^2-5x+6)}{(x-6)(x+1)}=\frac{3(x-2)(x-3)}{(x-6)(x+1)}\] ... Cannot be simplified...

OpenStudy (anonymous):

wat abt: -32a3b3 ---------- 48 a3 b3

OpenStudy (callisto):

Can you cross out the common factors first?

OpenStudy (anonymous):

i dont geret:/

OpenStudy (callisto):

\[\frac{-32a^3b^3}{48a^3b^3}\] What are the common factors of the denominator and numerator?

OpenStudy (anonymous):

A3 b3?

OpenStudy (callisto):

Yes, and?

OpenStudy (anonymous):

im lost lol

OpenStudy (callisto):

Hmm.. Let's cross out the common factors you've found first :| \[\frac{-32a^3b^3}{48a^3b^3} = \frac{-32}{48}\] How would you further simplify it?